OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2

OpenStudy (kropot72):

Do you want me to take you thru it step by step?

OpenStudy (anonymous):

Yes Please

OpenStudy (kropot72):

3x - 4y = -10 ...............(1) 5x + 8y = -2 ................(2) If we multiply both sides of equation (1) by 2 the coefficient of y will become -8. Then the resulting equation can be added to equation (2) and we will eliminate the y terms. Can you multiply both sides of equation (1) by 2?

OpenStudy (anonymous):

By 2

OpenStudy (kropot72):

2(3x - 4y) = 2(-10) Now you multiply out both sides. Can you do it?